(d) take half of r (e) double the magnitude of one of the charges (f) take half of one of the charges and double the distance between the two charges. Problem 2 Three charges, q,--1.51C, g2 SpiC, and g-21.c are placed as shown. where, there is a charge gn is at the origin, charge ga is at (4cm, 0cm), and charge (a is at (Ocm, 3cm). 0.3 0.4 О-У) 21.6 ス о- 16,845 Qyn 15.27 # AME
(A) Find the net electric force on charge q2 (Include Magnitude and Direction).
Answer #1


Summary of answer:-We use coulomb's law to find the electrostatic force due to q1 and q3 on charge q2.Now we write them in vector form so that on adding both the forces directly we will get vector form of net force.Now on taking the magnitude of resulting vector by analytical formula we get magnitude of net force on q2.To find the direction of net force with positive direction of x axis we use the formula arc tan(coefficient of j/coefficient of i).

(3)다(4) 2 るOm 0 23 change a Cat aetive Similar- 23 AB2 Fai 3 A 12 4 4k Onn3 = 17:201-1896 Net forca en 2 -う = _ as, 31 +17-22た1296の 三 ー8031-12-96J tude 803 ー) = 267-270